# MedCLIP-SAMv2: Towards Universal Text-Driven Medical Image Segmentation

### v3

@@ Abstract
Segmentation of anatomical structures and pathological regions in medical images is essential for modern clinical diagnosis, disease research, and treatment planning. While significant advancements have been made in deep learning-based segmentation techniques, many of these methods still suffer from limitations in data efficiency, generalizability, and interactivity. As a result, developing precise segmentation methods that require fewer labeled datasets remains a critical challenge in medical image analysis. Recently, the introduction of foundation models like CLIP and Segment-Anything-Model (SAM), with robust cross-domain representations, has paved the way for interactive and universal image segmentation. However, further exploration of these models for data-efficient segmentation in medical imaging is still needed and highly relevant. In this paper, we introduce MedCLIP-SAMv2, a novel framework that integrates the CLIP and SAM models to perform segmentation on clinical scans using text prompts, in both zero-shot and weakly supervised settings. Our approach includes fine-tuning the BiomedCLIP model with a new Decoupled Hard Negative Noise Contrastive Estimation (DHN-NCE) loss, and leveraging the Multi-modal Information Bottleneck (M2IB) to create visual prompts for generating segmentation masks from SAM in the zero-shot setting. We also investigate using zero-shot segmentation labels within a weakly supervised paradigm to enhance segmentation quality further. Extensive testing across four diverse segmentation tasks and medical imaging modalities (breast tumor ultrasound, brain tumor MRI, lung X-ray, and lung CT) demonstrates the high accuracy of our proposed framework. Our code is available at https://github.com/HealthX-Lab/MedCLIP-SAMv2.
{IEEEkeywords}
Text-driven Image segmentation, Vision-Language models, Foundation Models, Weakly Supervised Semantic Segmentation
@@ 1 Introduction
With the growing availability of radiological technologies, there is an increasing demand for precise and efficient medical image segmentation to support the study, diagnosis, and treatment of various medical conditions [1]. Deep learning (DL) techniques have emerged as state-of-the-art (SOTA) in this field; however, they face three key challenges that hinder their broader clinical use. First, the scarcity of large, well-annotated datasets presents a major obstacle to DL model development. Second, the lack of interactivity and interpretability undermines trust in these methods. Finally, most medical DL models are trained for specific tasks and contrasts/modalities, limiting their flexibility. While several self-supervised and weakly supervised approaches [2, 3, 4] have been introduced to improve training efficiency, and explainable AI (XAI) techniques (e.g., uncertainty estimation [5, 6] and saliency maps [7, 8]) are under active investigation, cross-domain generalization remains a major challenge.
@@
Recently, the introduction of foundation models, such as Contrastive Language-Image Pre-Training (CLIP) [9] and Segment Anything Model (SAM) [10] has paved the way for interactive and universal medical image segmentation. Several research groups have adapted CLIP and SAM for radiological tasks, including the development of BiomedCLIP [11] and MedSAM [12], which were pre-trained on vast amounts of biomedical data. However, further advances in parameter fine-tuning methods could enhance the performance of these models in radiology.
@@
As interest in SAM grows, to mitigate its reliance on drawing-based prompts (e.g., point and/or bounding box) for segmentation, which require prior clinical knowledge, recent methods have emerged to fine-tune SAM without prompts [13, 14], generate prompts via Class Activation Maps (CAM) from classification tasks [15, 16, 17], and refine its output using weak supervision [18, 19, 20]. While still in its early stages, the use of foundation models for interactive and universal medical image segmentation remains an important area for further exploration. To address these challenges, we introduce MedCLIP-SAMv2, a novel framework that integrates BiomedCLIP [11] and SAM [10] for text-prompt-based interactive and universal medical image segmentation, in both zero-shot and weakly supervised settings. The MedCLIP-SAMv2 framework extends and improves upon our original MedCLIP-SAM framework [21] from MICCAI 2024 by incorporating several updates:
@@
  • We explored different saliency map generation techniques, where we replaced gScoreCAM [22] with M2IB [23], which, when combined with our fine-tuning of BiomedCLIP [11], significantly improved zero-shot segmentation.
@@
  • We enhanced weakly supervised segmentation results from the previous framework by training nnUNet [24] using pseudo labels and providing uncertainty estimation via checkpoint ensembling [25].
@@
  • The validation was expanded by incorporating an additional Lung CT dataset, thereby covering four key radiology modalities—CT, MRI, ultrasound, and X-ray. This comprehensive testing demonstrates the framework’s versatility and robustness across diverse segmentation tasks.
@@
  • We investigated advanced text prompt engineering strategies by leveraging LLM reasoning and various ensembling methods, which significantly influence zero-shot segmentation performance.
@@
  • Significantly more extensive experiments were conducted for further validation, including testing different SAM backbones and visual prompts. We meticulously evaluate the necessity of each component in our framework and demonstrate the incremental contribution to the overall performance.
@@
The refined MedCLIP-SAMv2 framework is now more accurate, advancing further toward universal text-driven medical image segmentation. Our contributions are threefold: First, we propose a new CLIP training/fine-tuning approach called Decoupled Hard Negative Noise Contrastive Estimation (DHN-NCE). Second, we introduce a text-driven zero-shot medical segmentation method, combining CLIP and SAM for radiological tasks. Lastly, we explore a weakly-supervised strategy to further refine zero-shot segmentation results with uncertainty estimation. Our proposed framework is extensively validated across four distinct segmentation tasks and modalities, including breast tumor segmentation in ultrasound, brain tumor segmentation in MRI, and lung segmentation in chest X-ray and CT.
@@ 2.1 CLIP in Medical Domain
Several works have utilized CLIP for medical images and texts. Despite being trained on 400 million natural image-text pairs, CLIP’s performance suffers on medical tasks. For this reason, works like PubMedCLIP [26] suggested fine-tuning CLIP on a set of PubMed articles for the task of Medical Question-Answering; Zhang et al. [11] later showed PubMedCLIP’s poor performance on cross-modal retrieval tasks (worse than CLIP). On the other hand, MedCLIP [27] proposed a technique to utilize decoupled images and texts in the training process to augment data while Windsor et al. [28] explored different methods of enhancing the performance of vision-language models for medical domain tasks in a limited data setting. Alternatively, Wu et al. [29] proposed a method of enhancing the text in medical reports by simplifying the sentence complexity. Moreover, other works like [30] and [31] have utilized CLIP for the task of expert pathology detection and medical report generation. However, notably, almost all mentioned works [27, 28, 29, 30, 31] only utilized Chest X-ray data for their proposed methods. BiomedCLIP [11] is by far the most recent work for multi-modal medical data on a large scale, which was shown to be superior for cross-modal retrieval accuracy.
@@ 2.2 Weakly Supervised Semantic Segmentation
Many works have explored weakly supervised paradigms for segmenting distinct regions in natural images. CLIP-ES [32] proposed a purely text-driven approach to producing better pseudo-masks through CLIP’s class activation maps instead of training affinity networks, while SAMS [18] later extended the work by making use of the SAM model to produce coarse and fine seeds from image-level labels. Additionally, SG-WSSS [33] studied different methods, including scribbles, points, and bounding boxes to prompt SAM through CAM scores. However, these works may fail to translate well to medical scans, which have different characteristics than natural images. Novel CAM techniques specifically tailored for CLIP models like gScoreCAM [22] and M2IB [23] have emerged with SOTA results for generating multi-modal saliency maps. Specifically, gScoreCAM [22] utilized the top-K channel activations from the text and image encoder layers, leading to better-localized saliency maps. The more recent M2IB [23] reformulates the information bottleneck theory to multi-modal applications, where it was proven to outperform CAM-based, perturbation-based, and attention-based saliency mapping techniques. Additionally, M2IB also demonstrated its potential for medical image applications, where a fine-tuned CLIP model on Chest X-ray datasets was shown to properly highlight regions of abnormalities. Recently, Liu et al. [34] focus on improving interpretation of zero-shot medical image diagnosis through engineering relevant text prompts by integrating ChatGPT that outputs relevant descriptions of the radiological abnormality. However, the previous works don’t inspect improving medical segmentation through model training.
2.3 SAM for Medical Imaging Modalities
With the advent of SAM, a myriad of research has been dedicated to adapting it for medical imaging purposes. MedSAM [12] provided a large-scale fine-tuning of SAM on about 1 million medical image-mask pairs and demonstrated superior performance when it comes to multiple segmentation tasks. AutoSAM [35] offered a more efficient approach to fine-tuning SAM on medical images through training the prompt encoder and developing a lightweight deconvolution mask decoder for medical segmentation tasks. Cheng et al. [36] found that bounding boxes gave the best results when prompting SAM across 12 different medical segmentation tasks, and Huang et al. [20] proposed a pseudo-mask correction framework to enhance noisy labels generated from SAM for medical images that can be used for further fine-tuning. Finally, Gong et al. [37] replaced SAM’s mask decoder with a 3D convolutional neural network so that volumetric medical images can be supported.
@@ 3 Methods
An overview of the proposed MedCLIP-SAMv2 framework is presented in Fig. 1, which is organized into three distinct stages: BiomedCLIP fine-tuning employing our new DHN-NCE loss, zero-shot segmentation guided by text-prompts, and weakly supervised segmentation for potential label refinement.
@@
3.1 Efficient BiomedCLIP Fine-Tuning with the DHN-NCE loss
@@
Typically, a CLIP model is trained on large datasets consisting of images and their corresponding text descriptions. The model utilizes both an image encoder and a text encoder to extract features from images and texts, projecting them into vectors of the same dimension, denoted as $\mathbf{I}_{p,i}$ for images and $\mathbf{T}_{p,i}$ for texts. Through contrastive learning, the model learns a shared embedding space, where similar image-text pairs are positioned closer together while dissimilar pairs are pulled further apart. Although BiomedCLIP [11] was trained on medical charts/images and clinical texts, further fine-tuning can significantly enhance its performance on tasks specific to medical imaging. In traditional CLIP training with the InfoNCE loss [38], the negative-positive-coupling (NPC) effect [39] can reduce learning efficiency, especially with smaller batch sizes. Additionally, for medical images, distinguishing between subtle differences in cases within the same imaging category can be challenging. To address these issues, we propose the Decoupled Hard Negative Noise Contrastive Estimation (DHN-NCE) loss, which 1) combines the InfoNCE loss [38] with hard negative sampling [40], emphasizing “close samples”, and 2) incorporates decoupled contrastive learning [39] by removing the positive term in the denominator, allowing for smaller batch sizes. The resulting DHN-NCE loss function, $\mathcal{L}_{DHN-NCE}$, employs weighting functions ($\mathcal{W}_{\mathbf{I}_{p,i}\mathbf{T}_{p,j}}^{v\rightarrow t}$, $\mathcal{W}_{\mathbf{T}_{p,i}\mathbf{I}_{p,j}}^{t\rightarrow v}$) to increase the penalty for negative samples that are close to the anchor, using image-to-text and text-to-image hardness parameters $\beta_{1},\beta_{2}\geq 0$. In this context, $t\rightarrow v$ refers to text-to-image, while $v\rightarrow t$ indicates image-to-text.
@@
where $B$ is the batch size, $\tau$ is the temperature parameter, and the hardness weighting formulas are as follows:
@@
3.2 Zero-shot and Weakly Supervised Medical Image Segmentation
In this stage, we utilize the fine-tuned BiomedCLIP with the updated parameters $\boldsymbol{\theta}=\{\boldsymbol{\theta}_{\text{img}},\boldsymbol{\theta}_{\text{text}}\}$ as the backbone model for feature extraction from both images and text prompts. The core segmentation process relies on the Multi-modal Information Bottleneck (M2IB) technique, which generates visual saliency maps of the target regions by associating text prompts with image regions.
The zero-shot segmentation pipeline can be described mathematically as follows:

### v4

@@ Abstract
Segmentation of anatomical structures and pathologies in medical images is essential for modern disease diagnosis, clinical research, and treatment planning. While significant advancements have been made in deep learning-based segmentation techniques, many of these methods still suffer from limitations in data efficiency, generalizability, and interactivity. As a result, developing robust segmentation methods that require fewer labeled datasets remains a critical challenge in medical image analysis. Recently, the introduction of foundation models like CLIP and Segment-Anything-Model (SAM), with robust cross-domain representations, has paved the way for interactive and universal image segmentation. However, further exploration of these models for data-efficient segmentation in medical imaging is an active field of research. In this paper, we introduce MedCLIP-SAMv2, a novel framework that integrates the CLIP and SAM models to perform segmentation on clinical scans using text prompts, in both zero-shot and weakly supervised settings. Our approach includes fine-tuning the BiomedCLIP model with a new Decoupled Hard Negative Noise Contrastive Estimation (DHN-NCE) loss, and leveraging the Multi-modal Information Bottleneck (M2IB) to create visual prompts for generating segmentation masks with SAM in the zero-shot setting. We also investigate using zero-shot segmentation labels in a weakly supervised paradigm to enhance segmentation quality further. Extensive validation across four diverse segmentation tasks and medical imaging modalities (breast tumor ultrasound, brain tumor MRI, lung X-ray, and lung CT) demonstrates the high accuracy of our proposed framework. Our code is available at https://github.com/HealthX-Lab/MedCLIP-SAMv2.
keywords:
Text-driven Image segmentation , Vision-Language models , Foundation Models , Weakly Supervised Segmentation
journal: Medical Image Analysis
[…]
organization=Department of Electrical and Computer Engineering, Concordia University,city=Montreal, state=Quebec, country=Canada
[…]
organization=Department of Computer Science and Software Engineering, Concordia University,city=Montreal, state=Quebec, country=Canada
@@ 1 Introduction
With the growing availability of radiological technologies, there is an increasing demand for precise and efficient medical image segmentation to support the study, diagnosis, and treatment of various medical conditions (Siuly and Zhang, 2016). Deep learning (DL) techniques have emerged as state-of-the-art (SOTA) in this field; however, they face three key challenges that hinder their broader clinical adoption. First, the scarcity of large, well-annotated datasets presents a major obstacle to DL model development. Second, the lack of interactivity and interpretability undermines trust in these methods. Finally, most medical DL models are trained for specific tasks and contrasts/modalities, limiting their flexibility. While several self-supervised and weakly supervised approaches (Baevski et al., 2023; Chen et al., 2020; Taleb et al., 2021) have been introduced to improve training efficiency, and explainable AI (XAI) techniques, including uncertainty estimation (Loquercio et al., 2020; Liu et al., 2020) and saliency maps (Arun et al., 2021; Bae et al., 2020) are under active investigation, cross-domain generalization remains a major challenge.
@@
Recently, the introduction of foundation models, such as Contrastive Language-Image Pre-Training (CLIP) (Radford et al., 2021) and Segment Anything Model (SAM) (Kirillov et al., 2023) have paved the way for interactive and universal medical image segmentation. Several research groups have adapted CLIP and SAM for radiological tasks, including the development of BiomedCLIP (Zhang et al., 2023) and MedSAM (Ma and Wang, 2023), which were pre-trained on vast amounts of biomedical data. However, further advances in parameter fine-tuning methods could enhance the performance of these models in radiology.
Although CLIP training primarily operates at a global level for image-text mapping, research (Fu et al., 2024) has revealed that these models can encode rich feature representations of images. This allows us to establish the relationship between global textual information and local visual features (Zhou et al., 2022; Rao et al., 2022), which can be exploited for efficient zero-shot medical image segmentation, enabling broader applicability even in data-scarce settings, as we explored for the first time in our MICCAI 2024 paper (Koleilat et al., 2024b). The complex and nuanced nature of medical terminology, combined with the subtle and intricate variations inherent in medical images, introduces unique challenges that are less pronounced in natural images. While adapting CLIP to the medical image domain may seem attractive, it is non-trivial and requires substantial ground truth labels to fine-tune the model effectively, especially for segmentation downstream tasks (Poudel et al., 2023). The lack of large, high-quality annotated datasets in medical imaging further exacerbates this challenge. This calls for biomedical domain-specific CLIP models, such as BiomedCLIP (Zhang et al., 2023) and effective fine-tuning loss functions based on these domain-specific CLIP models to establish more effective cross-modal learning in radiological applications, such as pathology localization, segmentation, and diagnosis. We continue to explore these in this paper for MedSAM-CLIPv2.
@@
On the other hand, as interest in SAM grows, to mitigate its reliance on visual prompts (e.g., point and/or bounding box) for segmentation, which require prior clinical knowledge, recent methods have emerged to fine-tune SAM without these prompts (Chen et al., 2024; Hu et al., 2023), generate prompts via Class Activation Maps (CAM) from classification tasks (Li et al., 2024, 2023; Liu and Huang, 2024), and refine its output using weak supervision (Yang and Gong, 2023; Chen et al., 2023; Huang et al., 2023). While still in its early stages, the use of foundation models for interactive and universal medical image segmentation remains an important area for further exploration. Recently, to address these challenges, we introduced MedCLIP-SAM in MICCAI2024 (Koleilat et al., 2024b), which leverages BiomedCLIP (Zhang et al., 2023) to generate text-based box prompts for SAM (Kirillov et al., 2023) towards interactive and universal medical image segmentation, in both zero-shot and weakly supervised settings. Following the preliminary success, further improvement and exploration of the framework are necessary to further elevate the performance and gain deeper insights into the CLIP and SAM foundation models for medical imaging applications. As a result, in this paper, we propose MedCLIP-SAMv2, a novel technique that further evolves and significantly improves upon our original MedCLIP-SAM framework for zero-shot and weakly supervised medical image segmentation (Koleilat et al., 2024b). Specifically, the prominent upgrades for the newly proposed MedCLIP-SAMv2 framework from the original method include:
  • 1.
@@
    We investigated different saliency map generation techniques for CLIP models, where we replaced gScoreCAM (Chen et al., 2022) with M2IB (Wang et al., 2024), which, when combined with our fine-tuning of BiomedCLIP (Zhang et al., 2023), significantly improved zero-shot segmentation accuracy.
  • 2.
@@
    We enhanced weakly supervised segmentation results and interpretability from the previous framework by training nnUNet (Isensee et al., 2021) using pseudo-labels while providing uncertainty estimation via checkpoint ensembling (Zhao et al., 2022).
  • 3.
@@
    The validation was expanded by incorporating an additional Lung CT dataset, thereby covering four key radiological modalities — CT, MRI, ultrasound, and X-ray. This comprehensive testing further demonstrates the framework’s versatility and robustness across diverse segmentation tasks.
  • 4.
@@
    We investigated and optimized advanced text prompt engineering strategies by leveraging large language model (LLM) reasoning and various ensembling methods, which are shown to significantly boost zero-shot segmentation performance.
  • 5.
@@
    Significantly more extensive experiments were conducted for further validation of our framework’s design components, including testing different SAM backbones and visual prompt types. We meticulously evaluate the necessity of each component in our framework and demonstrate their individual contribution to the overall performance enhancement.
@@
The newly proposed MedCLIP-SAMv2 framework is more accurate, advancing further toward universal text-driven medical image segmentation with an increase of 13.07% and 11.21% in Dice score for zero-shot and weakly supervised paradigms, respectively. Our main contributions are threefold: First, we propose a new CLIP training/fine-tuning loss function called Decoupled Hard Negative Noise Contrastive Estimation (DHN-NCE). Second, we introduce a text-driven zero-shot medical segmentation method, combining CLIP and SAM for radiological tasks. Lastly, we explore a weakly-supervised strategy to further refine zero-shot segmentation results with uncertainty estimation. Our proposed framework is extensively validated across four distinct segmentation tasks and modalities, including breast tumor segmentation in ultrasound, brain tumor segmentation in MRI, and lung segmentation in chest X-ray and CT.
@@ 2.1 CLIP in Medical Domain
Several works have utilized CLIP for medical images and texts. Despite being trained on 400 million natural image-text pairs, CLIP’s performance suffers on medical tasks. For this reason, works like PubMedCLIP (Eslami et al., 2023) suggested fine-tuning CLIP on a set of PubMed articles for the task of Medical Question-Answering; Zhang et al. (Zhang et al., 2023) later showed PubMedCLIP’s poor performance on cross-modal retrieval tasks (worse than CLIP). On the other hand, MedCLIP (Wang et al., 2022) proposed a technique to utilize decoupled images and texts in training to augment data while Windsor et al. (Windsor et al., 2023) explored different methods of enhancing the performance of vision-language models for medical domain tasks in a limited data setting. Alternatively, Wu et al. (Wu et al., 2023a) proposed a method of enhancing the text in medical reports by simplifying the sentence complexity. Moreover, other works like (Keicher et al., 2023) and (Tiu et al., 2022) have utilized CLIP for the task of pathology detection and medical report generation. However, notably, almost all mentioned works (Wang et al., 2022; Windsor et al., 2023; Wu et al., 2023a; Keicher et al., 2023; Tiu et al., 2022) only utilized Chest X-ray data for their proposed methods. BiomedCLIP (Zhang et al., 2023) is by far the most recent work for multi-modal medical data on a large scale, which was shown to be superior for cross-modal retrieval accuracy. Notable studies (Koleilat et al., 2024a; Poudel et al., 2023) have investigated the transfer capabilities of BiomedCLIP in downstream tasks such as classification and segmentation. However, its adaptability remains largely unexplored compared to the extensive body of CLIP literature. To the best of our knowledge, our work is the first to explore the potential of BiomedCLIP in zero-shot segmentation tasks, paving the way for more efficient usage in medical imaging.
@@ 2.2 Weakly Supervised Semantic Segmentation
To mitigate the shortage of well-labeled datasets for medical image segmentation, many works have explored weakly supervised paradigms for segmenting distinct regions in natural images with CLIP-like models. CLIP-ES (Lin et al., 2023b) proposed a purely text-driven approach to producing better pseudo-masks through CLIP’s class activation maps instead of training affinity networks, while SAMS (Yang and Gong, 2023) later extended the work by making use of the SAM model to produce coarse and fine seeds from image-level labels. Additionally, SG-WSSS (Jiang and Yang, 2023) studied different visual prompting methods, including scribbles, points, and bounding boxes to prompt SAM through CAM scores. However, these works may fail to translate well to medical scans, which have different characteristics than natural images. Novel CAM techniques specifically tailored for CLIP models like gScoreCAM (Chen et al., 2022) and M2IB (Wang et al., 2024) have emerged with SOTA performance for generating multi-modal saliency maps. Specifically, gScoreCAM (Chen et al., 2022) utilized the top-K channel activations from the text and image encoder layers, leading to better-localized saliency maps. The more recent M2IB (Wang et al., 2024) reformulates the information bottleneck theory to multi-modal applications, where it was proven to outperform CAM-based, perturbation-based, and attention-based saliency mapping techniques. Additionally, M2IB also demonstrated its potential for medical image applications, where a fine-tuned CLIP model on Chest X-ray datasets was shown to properly highlight regions of abnormalities (Wang et al., 2024). Recently, Liu et al. (Liu et al., 2023) focus on improving interpretation of zero-shot medical image diagnosis through engineering relevant text prompts by integrating ChatGPT that outputs relevant descriptions of the radiological abnormality. However, these previous works don’t inspect improving medical segmentation through model training.
2.3 SAM for Medical Imaging Segmentation
With the advent of SAM, a foundation model for image segmentation that enables zero-shot generalization through a promptable architecture consisting of a powerful image encoder, a flexible prompt encoder, and a lightweight mask decoder, a myriad of research has been dedicated to adapting it for medical imaging purposes. MedSAM (Ma and Wang, 2023) provided a large-scale fine-tuning of SAM on about 1 million medical image-mask pairs and demonstrated superior performance when it comes to multiple segmentation tasks. AutoSAM (Shaharabany et al., 2023) offered a more efficient approach to fine-tuning SAM on medical images through training the prompt encoder and developing a lightweight deconvolution mask decoder for medical segmentation tasks. Cheng et al. (Cheng et al., 2023) found that bounding boxes gave the best results when prompting SAM across 12 different medical segmentation tasks, and Huang et al. (Huang et al., 2023) proposed a pseudo-mask correction framework to enhance noisy labels generated from SAM for medical images that can be used for further fine-tuning. Finally, Gong et al. (Gong et al., 2023) replaced SAM’s mask decoder with a 3D convolutional neural network so that volumetric medical images can be supported.
@@ 3 Methods
A full overview of the proposed MedCLIP-SAMv2 framework is presented in Fig. 3, which is organized into three distinct stages: 1) BiomedCLIP fine-tuning employing our new DHN-NCE loss, 2) zero-shot segmentation guided by text-prompts, and 3) weakly supervised segmentation for potential label refinement. We additionally showcase a summary of the main components of the framework in Fig. 1 for the readers’ easy reference.
Figure 1: A general overview of the essential components.
Figure 2: Comparison of the standard CLIP loss, which applies uniform penalties to all examples regardless of difficulty, with our DHN-NCE loss, which prioritizes harder examples. The DHN-NCE loss enhances the differentiation of medical cases by appropriately penalizing close negatives through adaptive weighting formulas. Green outline represents the anchor example while the red outline represents the negative examples.
@@
3.1 Efficient DHN-NCE Fine-tuning
@@
CLIP-like models are commonly trained on extensive datasets consisting of images paired with their corresponding textual descriptions. These models employ an image encoder and a text encoder to extract features, representing them as vectors in a shared dimensional space [^1] It is important to note that CLIP utilizes the global [CLS] tokens of the final vision and text encoder layers before projection to the shared embedding space: $\mathbf{I}_{p,i}$ for images and $\mathbf{T}_{p,i}$ for texts. Through the mechanism of contrastive learning, CLIP aligns semantically related image-text pairs by minimizing their distance in the embedding space while maximizing the separation of unrelated pairs. This shared embedding framework facilitates a cohesive understanding of multimodal data. Although BiomedCLIP (Zhang et al., 2023) was trained on medical charts/images and clinical texts, further fine-tuning can significantly enhance its performance on tasks specific to medical imaging. In traditional CLIP training with the InfoNCE loss (Oord et al., 2018), the negative-positive-coupling (NPC) effect (Yeh et al., 2022) can reduce learning efficiency, especially with smaller batch sizes. Additionally, for medical images, distinguishing between subtle differences in cases within the same imaging category can be challenging. To address these issues, we propose the Decoupled Hard Negative Noise Contrastive Estimation (DHN-NCE) loss, which 1) combines the InfoNCE loss (Oord et al., 2018) with hard negative sampling (Robinson et al., 2021), emphasizing “close samples”, and 2) incorporates decoupled contrastive learning (Yeh et al., 2022) by removing the positive term in the denominator, allowing for smaller batch sizes.
Original InfoNCE Loss: The standard InfoNCE loss for contrastive learning is formulated as follows:
$\mathcal{L}_{\text{InfoNCE}}=-\sum_{i=1}^{B}\log\frac{\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{i}^{+}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j}/\tau)}$ (1)
where $B$ is the batch size, $\mathbf{z}_{i}$ represents the feature embedding of the anchor sample, $\mathbf{z}_{i}^{+}$ is the positive pair for $\mathbf{z}_{i}$, $\tau$ is the temperature parameter, and $B$ is the batch size.
InfoNCE for Vision-Language Matching: To get a vision-language contrastive loss, we replace generic embeddings with image ($\mathbf{I}_{p,i}$) and text ($\mathbf{T}_{p,i}$) embeddings. In this context, $t\rightarrow v$ refers to text-to-image, while $v\rightarrow t$ indicates image-to-text:
$\mathcal{L}^{v\rightarrow t}=-\sum_{i=1}^{B}\log\frac{\exp(\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,i}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,j}/\tau)}$ (2)
$\mathcal{L}^{t\rightarrow v}=-\sum_{i=1}^{B}\log\frac{\exp(\mathbf{T}_{p,i}^{\top}\mathbf{I}_{p,i}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{T}_{p,i}^{\top}\mathbf{I}_{p,j}/\tau)}$ (3)
Decoupling Positives and Negatives: Expanding the logarithm and separating terms of Eq (2) and (3), we obtain:
$\mathcal{L}^{v\rightarrow t}=-\sum_{i=1}^{B}\left[\frac{\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,i}}{\tau}-\log\sum_{j=1}^{B}\exp(\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,j}/\tau)\right]$ (4)
$\mathcal{L}^{t\rightarrow v}=-\sum_{i=1}^{B}\left[\frac{\mathbf{T}_{p,i}^{\top}\mathbf{I}_{p,i}}{\tau}-\log\sum_{j=1}^{B}\exp(\mathbf{T}_{p,i}^{\top}\mathbf{I}_{p,j}/\tau)\right]$ (5)
Since the summation in the denominators of Eq (2) and (3) includes both the positive pair ($j=i$) and the negatives ($j\neq i$), we split it as:
$\sum_{j=1}^{B}\exp(\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,j}/\tau)=\exp(\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,i}/\tau)+\sum_{j\neq i}\exp(\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,j}/\tau)$ (6)
Following the positive-negative decoupling approach in (Yeh et al., 2022), we remove the positive pair and obtain the decoupled vision-language contrastive loss:
$\mathcal{L}^{v\rightarrow t}=-\sum_{i=1}^{B}\frac{\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,i}}{\tau}+\sum_{i=1}^{B}\log\left(\sum_{j\neq i}e^{\mathbf{I}_{p,i}^{\top}\mathbf{T}_{p,j}/\tau}\right)$ (7)
$\mathcal{L}^{t\rightarrow v}=-\sum_{i=1}^{B}\frac{\mathbf{T}_{p,i}^{\top}\mathbf{I}_{p,i}}{\tau}+\sum_{i=1}^{B}\log\left(\sum_{j\neq i}e^{\mathbf{T}_{p,i}^{\top}\mathbf{I}_{p,j}/\tau}\right)$ (8)
Applying Hardness Weights: The resulting DHN-NCE loss function, $\mathcal{L}_{DHN-NCE}$, employs weighting functions ($\mathcal{W}_{\mathbf{I}_{p,i}\mathbf{T}_{p,j}}^{v\rightarrow t}$, $\mathcal{W}_{\mathbf{T}_{p,i}\mathbf{I}_{p,j}}^{t\rightarrow v}$) to increase the penalty for negative samples that are close to the anchor, using image-to-text and text-to-image hardness parameters $\beta_{1},\beta_{2}\geq 0$.
@@
where the hardness weighting formulas are as follows:
@@
The weighting functions leverage exponential scaling to amplify the contributions of hard negatives—those with higher similarity scores—while suppressing easier negatives, ensuring the total weight distribution prioritizes these challenging cases (see Fig 2). By decoupling the positive term from the denominator, DHN-NCE prevents easy positives from diminishing the gradients associated with hard negatives. This mechanism sharpens the model’s focus on refining distinctions for harder cases, enabling more efficient training even with small batch sizes. Such properties make DHN-NCE particularly suited for medical imaging tasks with limited data availability and subtle feature variations.
3.2 Zero-shot Medical Image Segmentation
In this stage, we utilize the fine-tuned BiomedCLIP with the updated parameters $\boldsymbol{\theta}=\{\boldsymbol{\theta}_{\text{img}},\boldsymbol{\theta}_{\text{text}}\}$ as the backbone model for feature extraction from both images and text prompts. The core segmentation process relies on the Multi-modal Information Bottleneck (M2IB) technique (Wang et al., 2024), which generates visual saliency maps of the target regions by associating text prompts with image regions.
The zero-shot segmentation pipeline can be described as follows: